for the number 7,0,2,3,5 what is the maximum value and minimum value. what is their different value?
Answer by RAY100(1637)

arranging in order,,,0,2,3,5,7
min =0,,max=7,,range =(7-0)=7
,
looking at the problem another way
min value = 02,357=2,357
max value = 75,320
difference is =75320-2357=72,963
,,
please review problem set to decide which was required